OpenStudy (anonymous):

integrate (1/x^2) cos^2 (1/x) dx

OpenStudy (anonymous):

\(u=\frac{1}{x}\) is a start, and you get \(du=-\frac{dx}{x^2}\) so you are in luck

OpenStudy (anonymous):

okok. get it

OpenStudy (anonymous):

then how to integrate the cos^2??

OpenStudy (anonymous):

cos^2 u = (cos2u +1)/2 = (cos2u)/2 + 1/2 <- int this

OpenStudy (anonymous):